OpenStudy (anonymous):

Is this correct ? Find the sum of 2v6 and v24 * v stands for sqrt i got 2v30

OpenStudy (hoblos):

2√6 + √24 = 2√6 + √(4*6) = 2√6 +2√6 = 4√6

OpenStudy (anonymous):

\[\sqrt{24}=\sqrt{6*4}=\sqrt{6}*\sqrt{4}=\sqrt{6}*2=2 \sqrt{6}\]

OpenStudy (anonymous):

\[2\sqrt{6}+\sqrt{24}=2\sqrt{6}+2\sqrt{6}=4\sqrt{6}\]

OpenStudy (anonymous):

I guess I need to learn to type faster....lol